### v1

Abstract
Natural products, as metabolites from microorganisms, animals, or plants, exhibit diverse biological activities, making them crucial for drug discovery. Nowadays, existing deep learning methods for natural products research primarily rely on supervised learning approaches designed for specific downstream tasks. However, such one-model-for-a-task paradigm often lacks generalizability and leaves significant room for performance improvement. Additionally, existing molecular characterization methods are not well-suited for the unique tasks associated with natural products. To address these limitations, we have pre-trained a foundation model for natural products based on their unique properties. Our approach employs a novel pretraining strategy that is especially tailored to natural products. By incorporating contrastive learning and masked graph learning objectives, we emphasize evolutional information from molecular scaffolds while capturing side-chain information. Our framework achieves state-of-the-art (SOTA) results in various downstream tasks related to natural product mining and drug discovery. We first compare taxonomy classification with synthesized molecule-focused baselines to demonstrate that current models are inadequate for understanding natural synthesis. Furthermore, by diving into a fine-grained analysis at both the gene and microbial levels, NaFM demonstrates the ability to capture evolutionary information. Eventually, our method is experimented with virtual screening, illustrating informative natural product representations that can lead to more effective identification of potential drug candidates.
1 Main
Natural products are the aggregate of metabolic or secondary metabolic products produced by bacteria, fungi, animals, plants, and etc. The diversity among natural products is immense, encompassing hundreds of thousands of different structures [1]. As metabolites, natural products are more likely to exhibit bioactivity, thereby enhancing their potential to be directly developed into pharmaceuticals or structurally modified to serve as therapeutic agents, particularly in the development of anti-cancer and anti-infective agents [2, 3, 4]. Although finding a compound from nature that can be directly developed into a drug or precursor is highly valuable, it also entails significant time consumption and trial-and-error costs [5]. Traditional bioactivity-guided natural product discovery begins with the extraction of various natural products from the metabolites of microorganisms using solvents of different polarities. The crude extracts are then subjected to successive bioactivity-guided fractionation until pure bioactive compounds are obtained [6]. However, this process often requires identifying the structural classification of natural products and determining the specific microbial species responsible for producing the bioactive compounds to avoid unnecessary cultivation costs. Additionally, when identifying bioactive compounds through phenotypic assays, a considerable amount of time and effort is typically required to determine the molecular targets affected by these compounds [7].
Nowadays, with the advent of large-scale computing and chemical databases [8, 9, 10], activities [11], biological origins [11, 12], biosynthetic gene clusters (BGCs) [13], and specialized databases like the marine natural products database [14, 15, 16], natural product based drug design have move on to the next era. In this context, machine learning methods have been developed to mine valuable insights within the natural products data [17, 18, 19, 20, 21, 22]. For instance, Xu et al. [23] utilized deep learning techniques for the taxonomical classification of natural products, while Stokes et al. [24] employed deep neural networks to mine potential antibiotics. Building upon traditional fingerprint, e.g., ECFP, these methods could achieve satisfying performance but always suffer when applied to out-of-distribution test data or less-related downstream tasks. Tools commonly used for natural product classification, such as ClassyFire [25] and NPClassifier [26], are trained purely through supervised learning. However, due to the small amount of labeled data and data bias, they suffer from inaccurate classification results and poor generalization ability. Nevertheless, they are still widely used by multiple natural product-related databases as classification tools [9, 16, 12].
To address the aforementioned challenges, one effective approach is to pre-train a model on a large-scale unlabeled dataset. [27]. Molecular pre-training has a long history. Early approaches often adopt natural language processing techniques, representing molecules as SMILES [28] and utilizing RNNs [29] or Transformers [30] to learn molecular representations via masked language modeling [31, 32]. Subsequently, many studies have demonstrated that encoding molecules with graph neural networks [33, 34] and designing unique pre-training methods specifically for molecular graphs often yields superior performance [35, 36, 37]. More recently, research has shown that integrating 2D graph pre-training with various modalities of prior information about molecules, such as 3D structural data, semantic information, or force field data, can achieve even better performance [38, 39, 40, 41].
Despite extensive research on molecular pre-training, few attempts have been made to develop a specific foundation model for natural products. Natural products differ substantially from synthetic molecules, exhibiting greater structural diversity and complexity. They often possess unique scaffolds, stereochemistry, and functional groups rarely found in synthetic libraries [42]. Their intricate three-dimensional structures often make traditional methods struggle to capture 3D information or force field data, rendering such data less effective for representing natural products. Due to these challenges, it requires an alternative approach to effectively modeling natural products. Fortunately, natural products possess a distinctive property that can be leveraged. Similar to the central dogma that links DNA, RNA, and proteins, a fundamental relationship exists among the biological source, scaffold architecture, and biochemical properties of natural products. Specifically, natural products derived from the same biological source often share similar genes and biosynthetic gene clusters (BGCs). These genes encode enzymes or protein substrates that manipulate specific metabolic pathways, frequently producing conserved scaffolds. For example, the mevalonate (MVA) pathway commonly generates terpenoid or steroid scaffolds. Studies have further demonstrated that natural products with identical scaffolds exhibit remarkably similar biochemical properties [43]. This inherent relationship spans upstream genetic clusters and biosynthesis of natural products, extending to downstream bioactivity as shown in Fig. 1(d). Therefore, when modeling natural product structures, conventional molecular graph learning methods that distributing learning objectives across entire molecules fall short. Instead, we should design a novel hierarchical pre-training paradigm that focuses on scaffold-centered learning while effectively capturing side-chain structures, ultimately structuring the entire representation space for natural products.
Technically, current pre-training strategies have their own limitations that needs to be overcome. For instance, in existing works applying contrastive learning tasks to molecular pre-training, strategies like masking or removing substructures of molecules are used as data augmentation [35, 38]. Even though a similar strategy works for images, the chemical structures are much more sensitive to substructural changes. Minor structural changes can lead to significant variations in a molecule’s chemical or biological activity; this phenomenon is known as an activity cliff [44, 45, 46]. It indicates that in contrastive learning, we cannot simply use molecules with minor structural changes as positive examples [47]. Moreover, different molecules may produce structurally similar or even identical negative examples, which contradicts the principles of contrastive learning. Moreover, the current masked graph modeling approach also has its limitations. Simply masking a single atom or chemical bond allows the model to easily obtain sufficient information from directly connected atoms and bonds. However, we aim for the model to consider a more global structure during training to predict the masked information effectively. Additionally, when provided with the topological information of a molecule, models can easily exploit shortcuts to achieve pre-training targets, potentially undermining the intended learning objectives. Research has shown that increasing the difficulty of pre-training is positively correlated with its effectiveness [48]. Therefore, we should adjust the masking strategy and introduce additional targets for predicting masked information, thereby significantly enriching the information that the model can learn.
Figure 1: (a) Overview of NaFM pre-training. After the natural product molecule undergoes scaffold-aware masking, both the masked and unmasked information are simultaneously input into a multi-layer Graph Isomorphism Network. The masked information is then processed through a projection head to predict the masked atom attributes, bond attributes, and topological information. Meanwhile, the masked and unmasked information passes through a pooling function and is used for scaffold-aware contrastive learning. (b) Details of scaffold-subgraph reconstruction. First, a subgraph is randomly selected from the scaffold, consisting of multiple atoms and chemical bonds. In the subgraph, both node and edge attributes are masked, and all nodes within the subgraph are fully connected (i.e., expanded into a fully connected graph), thereby masking the topological information. During the reconstruction process, the model needs to predict both node and edge attributes, while also distinguishing between real edges and those artificially added virtual edges. (c) Details of scaffold-aware contrastive learning. Different colors represent varying weights defined by scaffold similarity. (d) The "central dogma" of natural products. The biological source, biosynthetic gene clusters, biosynthetic pathways, and bioactivity of natural products are interconnected through the scaffold, which acts as a bridge linking these three key aspects.
In this work, we have pre-trained a foundation model for small-molecule natural products (NaFM), that seamlessly combines masked learning and contrastive learning shown in Fig. 1(a). For the masked learning task, we propose a novel masked learning approach that goes beyond masking only categorical features of atoms and bonds. In addition to these, we also mask the connectivity information between atoms, effectively concealing the topological structure of subgraphs. To reconstruct these subgraphs, the model must first identify the locations of the bonds and then determine their types. Inspired by the fact that scaffolds carry coarse-grained biosynthetic information, while side chains result from post-synthesis modifications, we introduce scaffold similarity as a weighting function in our contrastive learning pretraining. This soft-weight approach enables the model to differentiate between strong and weak negative examples based on scaffold similarities, while also integrating valuable information from the side chains. Moreover, in our contrastive learning setup, the positive examples consist of masked molecules, while the negative examples come from the remaining molecules and their masked variants. This dual setup facilitates mutual supervision between the methods, leading to more effective learning and improved performance.
By specifically designing pre-training tasks tailored for natural products, NaFM is well-suited for a variety of significant and popular downstream tasks related to natural products. Starting from the structure, we applied NaFM to natural product taxonomy classification, a popular task in the field [49, 25, 50]. Specifically, we fine-tuned the pre-trained model on the original training dataset provided by NPClassifier [26] and replicated the representation learning baselines on this dataset to compare the performance in natural product classification. We further evaluate the pretrained model’s capacity to capture biosynthesis in LOTUS database [10]. Compared to supervised classification models, the pre-trained model significantly outperformed others by clearly distinguishing metabolites from different biological sources in the representation space. Leveraging biosynthetic gene clusters (BGCs) from MiBIG database [13], we conducted unprecedented genomic mining through metabolite structures. NaFM was able to identify BGCs and protein families based on metabolite structures. These results suggest that the pre-trained model can learn implicit genomic information from molecular structures. Finally, we selected several popular human protein target activity data from the NPASS database [11] for regression experiments. The results demonstrated that NaFM outperforms the baseline in directly predicting target activity. In terms of activity screening, we performed a detailed docking case study focusing on acetylcholinesterase (AChE). The docking results indicated that the molecules selected by NaFM exhibited better AChE inhibition activity compared to the control group.
2 Results
2.1 Overiew of NaFM
NaFM is a pre-training framework specifically designed for natural products. This framework integrates two novel pre-training tasks: scaffold-subgraph reconstruction and scaffold-aware contrastive learning.
Fig. 1(b) shows the details of scaffold-subgraph reconstruction. Basically, we consider a molecular graph where each node corresponds to an atom and each edge denotes a chemical bond connecting the atoms. During graph masking, we exclusively employ a subgraph masking strategy, and in addition to masking atom and bond information, we introduce the masking of topological information. Many natural product scaffolds feature unique polycyclic structures, with atom connectivity being more intricate and complex compared to synthetic molecules. To address this, masking the topological information not only enriches the information learned by the model but also increases the challenge of reconstructing atom and bond information. To further focus the model’s attention on the scaffold, we apply subgraph mask specifically to the scaffold portion of the molecular graph.
In contrastive learning, transformed objects originating from the same data points are treated as positive samples, while those derived from different objects are treated as negative samples. Although this approach is effective in traditional data forms like images, its application to molecular data, particularly natural products, presents challenges. This complexity arises because substructures from the same molecule can exhibit significant differences after augmentation, while different molecules may share identical substructures. Our goal is to obtain pre-trained models capable of producing effective structural representations; however, potential mislabeling may lead to suboptimal model performance. Due to the inherent information in natural product scaffolds, such as biological origin, biosynthesis, classification, and bioactivity, we incorporate scaffold similarity into contrastive learning to guide the model in distinguishing between different negative samples. That is to say, representations are pushed away based on whether they belong to a similar biosynthesis pathway.
To seamlessly integrate the two training tasks, the masked graphs are used not only for the scaffold-subgraph reconstruction task but also serve as positive examples in contrastive learning. In practice, contrastive learning generates a $2N\times 2N$ logits matrix shown in Fig. 1(c), where the positive example for each complete molecule is its corresponding masked graph, and the negative examples are formed by other molecules and their masked graphs.
This approach fosters the learning of richer and more accurate molecular representations. A comprehensive explanation of our methods is provided in the subsequent sections.
During the pre-training phase, we use approximately 0.6 million unlabeled data samples from the COCONUT database [9]. For downstream tasks, the parameters of the GNN backbone are retained from pre-training, while the prediction head is randomly initialized. The entire model is then fine-tuned using supervised learning on a dataset specifically curated for downstream tasks. Further details of this process are provided in the Methods section.
2.2 Natural Product Taxonomy Classification
The taxonomy classification of natural products is a classical and significant task. A valid natural product classification system could be valuable for database mining. Taxonomy labels are applied to infer structure-activity relationships, enabling preliminary predictions of their potential biochemical activities.
We mainly evaluate the classification performance on the NPClassifier dataset [26], which includes seven biosynthetic pathways of natural products and detailed classifications (SuperClass and Class) based on their structures. After cleaning the dataset by removing labels and corresponding data that were too sparsely represented, we obtained approximately 77,000 entries, with labels spanning 7 types of biosynthetic pathways, 70 SuperClasses, and 563 Classes.
We first fine-tuned the model at the Class level. Multiple pre-trained molecular models were used as baselines, replicated, and fine-tuned on this task for comparative analysis. We also selected the widely used non-pretrained representation, Extended-Connectivity Fingerprint(ECFP), as an additional baseline. When splitting the dataset into training, validation, and test sets, we ensured data balance by stratified splits based on the classes. During the comparison, we limited the number of molecules per class in the training set to simulate the classification performance of models under datasets of varying sizes. As shown in Table 1, it can be observed that under different testing dataset sizes, NaFM demonstrates significant performance improvements compared to other baselines, with even more pronounced enhancements when the dataset is smaller. The performance from graph neural network based methods, e.g. MolCLR, PretrainGNN, and D-MPNN express much more fine-tuning data size dependence as their performances drop drastically as we decreased the samples for fine-tuning. GraphMVP has achieved reasonable performance by incorporating implicit 3D information, while Mole-BERT benefits from its prior tokenizer training, which divides atoms of the same type into multiple distinct tokens. This approach increases the number of categories, thereby raising the difficulty of the masked atom modeling task. Overall results suggest that NaFM is the only pretraining model that outperforms rule-based fingerprint ECFP, which is the previous common choice for natural product classification.
Figure 2: The bar plots illustrate the models’ performance comparison of NaFM and NPClassifier on various superclass categories (a) and biosynthetic pathways (b) including Carbohydrates, Amino Acids and Peptides, Alkaloids, Terpenoids, Shikimates and Phenylpropanoids, Polyketides, and Fatty Acids in terms of AUPRC. The gray segments indicate the shared performance (minimum AUPRC) and the light blue/orange segments highlight the additional AURPC achieved by NaFM and NPClassifier beyond each other, respectively. Higher values indicate better model performance.
In addition to comparing NaFM with general molecular learning frameworks, we also benchmarked it against tools specifically designed for natural product classification. NPClassifier was chosen due to its open-source database, scientifically robust classification system, and its status as one of the most widely used methods for natural product classification [9, 16, 12]. We directly employed NPClassifier for comparison, fine-tuning the pre-trained model on the original NPClassifier dataset and conducting evaluations at both the Pathway and SuperClass levels. The results, presented in Fig. 2, show that NaFM significantly outperforms NPClassifier across all total biosynthetic pathways, as well as at the more granular SuperClass level. At the biosynthetic pathway level, our model outperforms NPClassifier by 24% in the recognition of amino acids and peptides. At the SuperClass level, our model outperforms 53 out of 71 superclasses, with significant improvements of over 20% in the identification of Carotenoids (C45), Tetramate alkaloids, Peptide alkaloids, and Anthranilic acid alkaloids. The gap in the recognition of $\gamma$-lactam-$\beta$-lactones can be attributed to the limited data for this category, with only 8 instances out of 78,000 data points, resulting in just one data point in the test set. These results suggest that NaFM can be further instrumentalized as the most accurate natural product classification tool to date for various natural product-related databases.
2.3 Biochemical significance of NaFM representation
Figure 3: Natural Products Atlas. The obtained representations are projected into two dimensions using TMAP [51] and visualized with Faerun [52]. For each biological source, a representative example is highlighted in a designated box, showcasing a characteristic scaffold that typifies compounds derived from that particular source.
In this section, we explore the biological information integrated within our pretraining models. For natural products that are challenging to synthesize, tracing the pathways of these compounds [53, 54, 23] facilitates combinatorial biosynthesis [55].
Given the strong interrelationship between molecular scaffolds, microbial species, and biosynthetic gene clusters, the specially designed pre-trained model performed exceptionally well in addressing both tasks. We fine-tune the model using the LOTUS dataset [10], which contains 130,000 data points with biological source labels. As illustrated in Fig. 4, NaFM effectively separates natural products from five major biological sources—animals, bacteria, chromista, fungi, and plants—within the representation space. Research on chromista is still emerging and has not kept pace with studies on other biological sources. Their unique environments and the difficulties in obtaining samples have resulted in few natural products being identified from them, with only about 1% of data coming from chromista. Nevertheless, despite the scarcity of training data, NaFM is still able to effectively differentiate natural products from chromista. For each biological source, we performed structural clustering and selected the representative scaffold corresponding to the cluster with the most data. This scaffold reflects the most characteristic structure of natural products from that source. The distinct differences in scaffolds across biological sources further highlight the strong relationship between the source and the scaffold structures. In contrast, when the same model architecture is applied without loading the pre-trained weights, the model is unable to distinguish between natural products from these sources. This result further underscores the advantages of NaFM over approaches that rely on direct supervised learning using labeled datasets.
Genetic studies of natural products have revealed that the synthesis of bioactive molecules is primarily governed by groups of co-localized genes, known as biosynthetic gene clusters (BGCs) [56, 57]. Traditional methods typically rely on analyzing the structure of BGCs to predict metabolite structures or to identify BGCs within genomes [21, 58, 59]. In contrast, reverse prediction of BGCs based on metabolite structures allows synthetic biologists to more precisely pinpoint the genomic regions responsible for synthesizing specific metabolites. This strategy not only enhances the accuracy of BGC selection for engineering purposes but also facilitates the large-scale production of these metabolites and streamlines the overall production process. To this end, we utilized information from over 2,000 bacterial and fungal BGCs and their corresponding metabolite structures from MIBiG [13], along with the protein family data for these BGCs from Pfam [60], to create a dataset containing metabolite structures and the corresponding BGC-associated protein families. Subsequently, we applied NaFM and various molecular representations to the task of predicting BGCs which is encoded by pfam based on metabolite structures. As shown in Fig. 3, we selected the 128 most frequent protein families in the dataset (detailed information on these protein families is shown in Supplementary Table S6) and analyzed the classification performance of these models across these protein family categories. It is evident that NaFM outperforms the other molecular representation methods overall, with a significant advantage over MolCLR, ECFP, and PretrainGNN.
Specifically, analyzing the blue blocks in the figure, the model shows significantly better classification performance on certain protein families, such as the Conserved Protein Domain Family (ID: 65 in the figure) and HATPase_c (ID: 115 in the figure), compared to other baselines. The Conserved Protein Domain Family includes protein domains that are widely conserved across different species and are involved in a variety of biological processes, indicating structural and functional similarity [61], while the HATPase_c family is associated with ATP hydrolysis, playing a key role in cellular processes like signal transduction and DNA repair [62]. Both of these families are considered conserved protein families, which typically exhibit high sequence and structural similarity. These results suggest that NaFM is capable of making effective predictions by learning the conserved features within these structures. In addition to analyzing the top 128 protein families, we also compared the average classification performance across all protein families. As shown in Supplementary Figure S3, NaFM significantly outperforms the other baselines in terms of both AUROC and AUPRC, with a smaller variance. Compared to ECFP, both metrics improved by more than 10%. In summary, NaFM demonstrates exceptional predictive performance at the genetic level, highlighting its potential as a powerful tool for BGC mining in the future.
Figure 4: The figure visualizes the AUROC gap between NaFM and various molecular representations. The results show the AUROC for the 128 most frequent protein families found in bacteria and fungi. The color in each block represents the magnitude of the AUROC gap, ranging from negative to positive, with colors transitioning from red to blue. Deeper colors indicate a greater absolute difference.
2.4 Natural product bio-activity prediction and virtual screening
Figure 5: (a) Docking scores of model ranked top 0.1% (approximately 600 molecules) of acetylcholinesterase inhibitors, along with 600 randomly selected molecules. The histogram and distribution of docking scores with acetylcholinesterase are shown in the figure. Orange represents the model-selected group, while blue represents the random group. The x-axis represents the docking score, and the y-axis represents frequency. A lower docking score (more left-shifted distribution) indicates higher inhibitory activity. Docking was performed using Schrödinger Maestro, with detailed settings provided in the Supplementary section 1.3. (b) A molecule with a high model score was selected for docking visualization. This molecule forms cation-$\pi$ interactions with the residues PHE-388, TYR-337, and TRP-86 in the acetylcholinesterase binding pocket (yellow dashed lines), and a hydrogen bond interaction with the residue SER-125 (red dashed line). (c) The first and second columns display several molecules with high model scores and good docking activity, along with their most similar positive counterparts encountered during the model fine-tuning process. The third column shows the maximum common substructure between these two molecules. (d) The enrichment factor results of NaFM and Glide-based molecular docking for activity screening across seven targets. The values are obtained through five-fold cross-validation, with the mean and variance reported. A higher enrichment factor indicates a stronger screening capability of the method.
Due to the difficulty in synthesizing or extracting natural products, there is significantly less activity data available for them compared to synthetic molecules [63]. Consequently, the knowledge learned through fine-tuning is limited, posing a challenge for the amount of information the model can acquire directly during pre-training. Although cell line data are more abundant, their complexity—stemming from multiple proteins, organelles, and signaling pathways—renders the structure-cytotoxicity relationship intricate. In contrast, human protein target data, with defined 3D structures and binding sites, exhibit clearer structure-activity relationships through specific molecular interactions (e.g., hydrogen bonds, hydrophobic and ionic bonds). Therefore, we focused solely on human protein targets for subsequent activity prediction and screening experiments.
To validate the activity prediction capability of NaFM, we selected human protein targets and HIV type-1 reverse transcriptase activity data from the NPASS natural product activity database, with activity data greater than 100 for each target. We then fine-tuned NaFM and various molecular representation baselines on this refined dataset and recorded the Root Mean Square Error(RMSE) for each method across different protein target activity prediction tasks. The results are shown in Table 2. As demonstrated, NaFM surpasses all baselines, with the exception of N-Gram on HIV type-1 reverse transcriptase, where NaFM performs marginally less effectively. However, it should be noted that the HIV reverse transcriptase target is quite different from human protein targets in terms of its structure and functional mechanisms, which can lead to some confusion in the model. In contrast, N-Gram, based on a statistical language model, is not affected by such structural complexities.
NaFM shows notable improvements over models like GraphMVP and PretrainGNN across various protein targets, such as PTP1B, AChE, COX-2, Tyrosinase, and CYP3A4. Specifically, NaFM outperforms these models in a clear and consistent manner, achieving lower RMSE values for most targets. For example, the RMSE for NaFM on COX-1 is 0.9286, compared to GraphMVP’s 3.0083 and PretrainGNN’s 2.6831. The improvements are particularly significant in the context of more complex and data-scarce protein targets, where the model shows superior generalization capability.
The reason for the relatively poor performance of GraphMVP can be attributed to its use of molecular geometry information during pre-training. Given the structural complexity of natural products, the 3D atomic coordinates and bond lengths/angles predicted by RDKit are often inaccurate. This introduction of erroneous geometry information during training can negatively impact the model’s performance, making it less effective at handling the unique complexities of natural product structures.
In addition to activity prediction, virtual screening is a traditional yet crucial method in drug discovery. It is aimed at identifying molecules from large compound libraries that are most likely to bind to a specific target and exhibit the desired biological activity, thereby facilitating the discovery of potential drug candidates [64]. Common virtual screening methods include molecular docking, which predicts interactions between amino acid residues in the receptor pocket and the ligand [65, 66, 67], as well as deep learning methods for binding-affinity prediction [68, 69, 70, 71]. While much research has been conducted on the virtual screening of synthetic small molecules, virtual screening of natural products holds additional significance. Due to the challenges in purifying natural products and the fact that their structures often contain multiple chiral centers, fused ring systems, and reactive functional groups (such as hydroxyl groups), their synthesis is highly complex and difficult. As a result, traditional high-throughput screening (HTS) methods often fall short in natural product research [72]. Virtual screening allows for the identification of bioactive natural product molecules that have not yet been systematically studied, potentially providing new insights.
In this context, we apply the rich representations learned by NaFM to the virtual screening of natural products. To evaluate the performance of the virtual screening, we selected acetylcholinesterase(AChE) as the target. AChE has a single active site, and its inhibitors are commonly used to treat neurodegenerative diseases such as Alzheimer’s, Parkinson’s, and Lewy body dementia [73]. We first use the AChE IC50 data from the NPASS database, categorizing compounds as active or inactive based on a micromolar cutoff. The model is then fine-tuned on the activity data, and the fine-tuned model is then used to score the activity of compounds in the COCONUT dataset. We conducted molecular docking using Schrodinger Maestro [65] on the top 0.1% of approximately 600 molecules with the highest scores, alongside a randomly selected set of 600 molecules.
After preprocessing the protein crystal structure and the two sets of ligands, we obtain the docking scores, as shown in Fig. 5(a). We observe that the distribution of docking scores for both groups closely approximates a normal distribution, indicating that the sample size is sufficient and the results are valid. Compared to the docking score distribution of the randomly selected control group, there is a significant difference in the distribution of docking scores for the molecules selected by the model. The model-selected group not only has a higher mean but also a smaller variance, with docking scores predominantly concentrated between 8 and 10, suggesting excellent performance in the model’s virtual screening. Furthermore, no molecules in the model-selected group have docking scores below 4, and most molecules have docking scores no lower than 6, indicating a very low number of false positives in the molecules recalled by the model, thereby reducing the additional experiment costs associated with incorrect recall. In contrast, the control group, selected randomly, exhibits a lower mean docking score, larger variance, and a considerable number of molecules with low docking scores. Furthermore, we selected one of the highest-scoring molecules from the model and performed docking visualization using PyMOL, as shown in Fig. 5(b). It can be observed that the molecule interacts with several amino acid residues in the protein pocket. We select high-confidence interactions for display. The amino cation (NH2+) in the molecule forms cation-$\pi$ interactions with the aromatic rings of PHE-388, TYR-337, and TRP-86, while another nitrogen atom at the tail of the molecule forms a hydrogen bond with SER-125.
However, in many cases, the goal is not only to find active molecules but also to identify novel active compounds with significantly different scaffolds or overall structures compared to known active molecules. To facilitate the discovery of novel scaffolds, we leverage NaFM’s learned representations to search for potential leads that are structurally distinct from known actives. As shown in Fig. 5(c), we present high-scoring molecules identified by both the model and docking, alongside the most structurally similar positive molecules learned during fine-tuning. The third column displays the maximum common substructures (MCS) between molecule pairs from both groups. It is evident that there are few recurring structural fragments, demonstrating NaFM’s ability to effectively identify novel and active molecules. This capability provides strong support for the discovery of new lead compounds or fragments from natural products.
In order to further demonstrate NaFM’s capability in activity screening across diverse human protein targets, we fine-tuned the model on activity data from seven key human protein targets and followed the previous virtual screening procedure, using enrichment factors as the evaluation metric. Given that Glide is a widely used and powerful docking and virtual screening tool, we selected Glide as a baseline to compare the screening performance of NaFM. As shown in Fig. 5(d), NaFM outperforms Glide in EF1%, EF5%, and EF10% across all targets except for COX-1, where its EF1% is slightly lower than that of Glide. Notably, for targets such as AChE, PTP-1B, 5-Lipoxygenase, and P-glycoprotein 1, Glide exhibits little to no enrichment capability, whereas NaFM demonstrates excellent performance. The suboptimal performance of Glide in these cases may be attributed to the challenges associated with fitting the three-dimensional structures of natural products. These results indicate that NaFM can serve as a powerful and efficient virtual screening tool for natural products, enabling the rapid identification of bioactive compounds.
3 Conclusion
In this paper, we have proposed NaFM, a foundation model for small molecule natural products. NaFM leverages the power of a large amount of unlabeled natural product structures and an efficient biological insight-inspired pertaining model. Our method has been evaluated across a range of tasks, including upstream genome mining, biological source identification, structural classification, metabolic pathway prediction, and activity prediction and screening. On most of the benchmarks, NaFM outperforms both traditional molecular fingerprints and deep-learning based pre-training frameworks. In the future, NaFM holds great potential as a foundation model for natural product based drug discovery and biosynthesis studies. Furthermore, it can serve as a tool-driven platform that streamlines the entire research workflow of natural products, integrating computational analysis, predictive modeling, and experimental validation to accelerate discoveries in the field.
4 Methods
4.1 Initial Representations of molecules
In our work, a molecule is represented as an attributed graph $G=(V,E)$, where atoms are nodes ($V$) and chemical bonds are edges ($E$). We adopt the standard feature representation commonly used in previous studies [35]. As detailed in Supplementary Table S1, the initial node features are constructed by summing the embeddings of attributes including atom types, chirality, and formal charges. Similarly, the initial bond features are obtained by summing the embeddings of bond types and bond directions.
$\text{Representation}_{atom/bond}=\sum_{i=0}^{F_{atom/bond}}\text{Embed}_{i}\left({\text{Attribute}_{i}}\right)$ (1)
where, $F_{*}$ is the number of atom or bond features. These initial features also serve as labels for the reconstruction task. We define the possible categories as the Cartesian product of all attribute sets, resulting in $9\times 4\times 5=180$ distinct atom categories ($M_{atom}$) and $4\times 3=12$ distinct bond categories ($M_{bond}$). In other words, each unique combination of atom attributes forms an atom category, and each unique pair of bond attributes defines a bond category.
4.2 Model Architecture
We adopt the traditional message-passing scheme introduced in GIN [74], with modifications to incorporate edge features as proposed in [75]. The node representation is updated as follows:
$\hat{h}^{l+1}_{i}=\text{MLP}\left(h^{l}_{i}+\sum_{j\in\mathcal{N}(i)}(h^{l}_{j}+e_{ij})\right),$ (2)
where $h^{l}_{i}$ is the representation of node $i$ in the $l$-th layer, $e_{ij}$ represents the edge feature between nodes $i$ and $j$, and $\mathcal{N}(i)$ denotes the neighbors of node $i$. To stabilize training, we apply layer normalization after each GIN block:
$h^{l+1}_{i}=\text{LayerNorm}(\hat{h}^{l+1}_{i})$ (3)
The final node representations, $h^{L}$, are aggregated using an average pooling operation to derive the graph-level representation.
$h_{\text{graph}}=\frac{1}{|V(G)|}\sum_{v\in G}h_{v}$ (4)
where $|V(G)|$ is the number of nodes in each graph. For the contrastive learning task, an additional MLP is used to project the graph representation into the latent space. For reconstruction, the final node representations are mapped to the corresponding atom categories, while the final edge representations are obtained by concatenating the node representations and applying an MLP to adapt them to the specific tasks. During fine-tuning on downstream tasks, we remove the contrastive learning and reconstruction head. The final graph representation is passed through a newly added MLP for classification or regression tasks.
4.2.1 Scaffold-aware Contrastive Learning
The issue of mislabeling in contrastive learning can be mitigated through the application of re-weighting techniques [76]. However, determining appropriate weights is a challenging task. Interestingly, the distinctive characteristics of natural products—specifically, the strong correlation between their scaffold structures and their properties—offer a promising solution for weight selection: scaffold similarity. To implement this, we compute the cosine similarity between the MACCS fingerprints [77] of their scaffolds. This similarity is then incorporated into the contrastive learning loss function for re-weighting, as detailed below:
$\mathcal{L}_{c}^{w}=\frac{1}{N}\sum_{m=1}^{N}\left[\ell\left(2m-1,2m\right)+\ell\left(2m,2m-1\right)\right],\ell_{i,j}=-\log\frac{s_{i,j}}{s_{i,j}+\sum_{k=1}^{2N}\mathbb{I}_{[k\neq i,j]}(1-w_{i,k})s_{i,k}}$ (5)
where $s_{i,j}=\exp\left(\text{sim}(h_{i},h_{j})/\tau\right)$ is the similarity in exponential space between the representations of $i$-th molecule and $j$-th molecule in a batch contained $N$ samples and $\mathbb{I}$ is the indicator function.
The weighted loss function can be interpreted as assigning lower weights to negative sample pairs from different molecules that exhibit more scaffold similarity. This allows the model to reduce its attention to these "false negatives", preventing it from being misled. As a result, the model achieves a more robust representation of molecular topology structures. Although reweighting techniques were introduced, the model tended to overfit the contrastive learning task as training progressed, leading to attempts to distinguish these "false negatives". To further address this issue, we developed a cosine loss weight scheduler as follows:
$\lambda_{c}^{i}=\frac{1}{2}\cos(1+\frac{i\pi}{N})$ (6)
where $i$ is the current epoch and $N$ is the total epochs. In this way, the model can mitigate overfitting on this task while ensuring sufficient training for other tasks.
4.2.2 Scaffold-subgraph Reconstruction
With the data-augmented samples, which naturally provide labels for mask modeling, the two frameworks were seamlessly integrated. These two approaches are complementary: contrastive learning is prone to overfitting, as contrastive models may find shortcuts through trivial representations. However, with the aid of mask modeling, the model goes beyond merely distinguishing between positive and negative sample pairs. It leverages information from functional groups (non-masked) to predict scaffold information (local). Simultaneously, contrastive learning enhances the model by helping it learn a global semantic representation. It is important to note that we define the task as scaffold-subgraph reconstruction, while discarding traditional approaches such as atomic masking and bond deletion. These tasks are considered too simplistic for the model and do not contribute meaningful knowledge. To ensure that the model focuses on scaffold modeling, we limit the masked portions exclusively to the atoms within the scaffold. In the reconstruction phase, we introduce a novel loss function that combines three classification tasks: atom type prediction, link prediction, and bond type prediction:
$\mathcal{L}_{r}$ $=\mathcal{L}_{atom}+\mathcal{L}_{link}+\mathcal{L}_{bond}$ (7)
$=-\frac{1}{N_{atom}}\sum_{i}\sum^{M_{atom}}_{c=1}y_{ic}\log(p_{atom}^{ic})-\frac{1}{N_{bond}}\sum_{j}\sum^{2}_{c=1}y_{jc}\log(p_{link}^{jc})-\frac{1}{N_{bond}}\sum_{j}\sum^{M_{bond}}_{c=1}y_{jc}\log(p_{bond}^{jc})$ (8)
where, $N_{*}$ is the number of atoms or bonds, $M_{*}$ is the number of atom or bond category. Here, $y_{ic}$ represents the indicator function, where the value is 1 if sample $i$ belongs to category $c$, and 0 otherwise. Additionally, $p^{ic}_{*}$ denotes the probability that the model predicts sample $i$ belongs to category $c$. More detailed information about atom or bond categories is shown in Methods section.
4.3 Scaffold Masking Algorithm
Scaffold masking is a critical step in our framework, as it plays a key role in both the augmented data of contrastive learning and the reconstruction task. The masking process is described in detail below.
As shown in Algorithm 1, initially, one scaffold atom is randomly selected, and its corresponding node is removed. The process continues by identifying the neighbors of the currently selected nodes, prioritizing those in the scaffold. These neighboring nodes are shuffled and iteratively removed until the desired number of nodes is reached or no further eligible neighbors exist. The output is a tensor containing the indices of the removed nodes. After that, all removed nodes are marked as masked, and all edges containing any of the removed nodes are also masked to form a augmented graph of contrastive learning and input/output pair of the reconstruction task.
Input: Graph $G$, Scaffold IDs $S$, Mask ratio $m$
Output: Modified graph $G_{\text{new}}$, Masked nodes $V_{\text{masked}}$
1
2Initialization:
3 $M\leftarrow\max(\min(\lceil|V(G)|\cdot m\rceil,|S|),2)$
4 $V_{\text{masked}}\leftarrow\emptyset$
5 $V_{\text{temp}}\leftarrow\text{Random choice}(S,1)$
6
7while $|V_{\text{masked}}|<M$ do
8       $V_{\text{neighbors}}\leftarrow\emptyset$
9       foreach $n\in V_{\text{temp}}$ do
10             $V_{\text{neighbors}}\leftarrow V_{\text{neighbors}}\cup\{i\in\text{Neighbors}(G,n)\mid i\notin V_{\text{temp}}\text{ and }i\in S\}$
11
12       end foreach
13      $V_{\text{neighbors}}\leftarrow\text{Shuffle}(V_{\text{neighbors}})$
14       foreach $n\in V_{\text{temp}}$ do
15             if $|V_{\text{masked}}|<M$ then
16                   $G\leftarrow\text{Remove node}(G,n)$
17                   $V_{\text{masked}}\leftarrow V_{\text{masked}}\cup\{n\}$
18
19             end if
20            else
21                   break
22             end if
23
24       end foreach
25      $V_{\text{temp}}\leftarrow\text{unique nodes in }V_{\text{neighbors}}$
26       if $V_{\text{temp}}=\emptyset$ then
27             break
28       end if
29
30 end while
return $V_{\text{masked}}$
Algorithm 1 Masked Subgraph Generation
4.4 Training Details
Pre-training
We implement a 6-layer Graph Isomorphism Network (GIN) [74] with layer normalization as the GNN backbone, using a hidden dimensionality of 512. An average pooling operation is applied as the readout function to extract a global molecular representation of 1024 dimensions. For the contrastive learning (CL) task, a single MLP with LeakyReLU activations projects the representation to 256 dimensions. For the reconstruction task, three MLPs with LeakyReLU activations map the node or edge representations to their respective output dimensions. All MLPs use a hidden dimensionality of 512. The masking ratio is set to 0.2.
We optimize the weighted CL and reconstruction loss using the AdamW optimizer with a weight decay of $10^{-5}$. The temperature for the CL loss $\tau$ is set to 0.1. A cosine decay schedule is applied to the scaling factor of the CL loss, decreasing from 1 to 0. The learning rate is set to $1\times 10^{-4}$ with a cosine decay schedule. The model is trained with a batch size of 256 for a total of 300 epochs. To avoid overfitting, we also set the dropout ratio to 0.1. A detailed hyper-parameter search space is provided in Supplementary Table S2.
Downstream Fine-tuning
For downstream task fine-tuning, we replace the original contrastive learning (CL) head and reconstruction head with a newly added, randomly initialized MLP on top of the GNN backbone. The softmax cross-entropy loss is used for classification tasks. For each task, we use the AdamW optimizer with a weight decay of $1\times 10^{-5}$ and a learning rate of $5\times 10^{-4}$, along with a cosine learning rate decay schedule. The model is fine-tuned for 300 epochs with a batch size of 512. The dropout ratio in the newly added MLP is set to 0.3. Early stopping is applied if the evaluation metric does not improve for 20 consecutive epochs. Thanks to the robust performance of the pre-training, we do not perform hyper-parameter tuning for the downstream tasks.
For all pre-trained baselines, we use the same settings as described above. For baselines trained from scratch, we follow their default configurations.
5 Tables
Table 1: The AUPRC results for natural product taxonomy classification compared to other baselines. AUPRC values were computed three times using different random seeds, and the averages were reported. Standard deviations are indicated following "±". Bold numbers indicate the best performance.
Training samples per Class 4 8 16 24 40 64
N-Gram [78] $44.72\pm 1.91$ $56.61\pm 0.66$ $66.36\pm 1.57$ $71.11\pm 0.77$ $73.34\pm 1.14$ $75.77\pm 0.54$
PretrainGNN [36] $44.83\pm 0.82$ $61.85\pm 0.78$ $75.76\pm 0.62$ $80.50\pm 0.17$ $85.31\pm 0.36$ $87.82\pm 0.13$
D-MPNN [79] $46.63\pm 0.23$ $60.88\pm 0.48$ $75.73\pm 0.12$ $80.96\pm 0.24$ $86.64\pm 0.44$ $89.23\pm 0.79$
MolCLR [35] $45.76\pm 1.98$ $65.80\pm 1.51$ $78.14\pm 1.08$ $83.20\pm 0.74$ $85.56\pm 0.11$ $88.22\pm 0.34$
Mole-BERT [37] $66.32\pm 1.23$ $73.39\pm 0.66$ $78.25\pm 0.45$ $80.83\pm 0.51$ $83.57\pm 0.59$ $85.69\pm 0.70$
ECFP [80] $69.17\pm 0.19$ $78.21\pm 0.79$ $83.82\pm 0.37$ $86.28\pm 0.49$ $88.52\pm 0.48$ $89.75\pm 0.45$
GraphMVP [38] $64.50\pm 0.74$ $78.41\pm 0.09$ $85.71\pm 0.43$ $87.88\pm 0.26$ $89.72\pm 0.45$ $91.07\pm 0.43$
NaFM $\mathbf{70.10\pm 0.92}$ $\mathbf{79.89\pm 0.07}$ $\mathbf{87.37\pm 1.51}$ $\mathbf{89.15\pm 0.22}$ $\mathbf{90.77\pm 0.26}$ $\mathbf{91.75\pm 0.47}$
Table 2: The RMSE results for natural product bioactivity regression compared to other baselines. RMSE values were computed three times using different random seeds, and the averages were reported. Standard deviations are indicated following "±". Bold numbers indicate the best performance.
Target Name PTP1B AChE COX-2 HIV type-1 RT Tyrosinase CYP3A4 MRP4 COX-1
Number of samples 612 341 190 186 186 178 177 140
N-Gram [78] $0.8493\pm 0.2623$ $1.3774\pm 0.3995$ $1.0022\pm 0.3192$ $\mathbf{1.0606\pm 0.1613}$ $0.7822\pm 0.1408$ $0.7846\pm 0.2037$ $0.2747\pm 0.0846$ $0.9529\pm 0.1643$
PretrainGNN [36] $0.8652\pm 0.2110$ $1.5462\pm 0.2265$ $1.7231\pm 0.3754$ $1.5210\pm 0.2606$ $1.6957\pm 0.2915$ $1.4157\pm 0.2923$ $1.6869\pm 0.2476$ $2.6831\pm 0.2056$
D-MPNN [79] $0.8577\pm 0.2165$ $1.2881\pm 0.1992$ $1.0272\pm 0.2468$ $1.1826\pm 0.1683$ $0.8206\pm 0.1891$ $0.6946\pm 0.1415$ $0.2399\pm 0.0603$ $0.9405\pm 0.1414$
MolCLR [35] $0.9912\pm 0.2467$ $1.1507\pm 0.1683$ $0.9399\pm 0.1705$ $1.2649\pm 0.1353$ $0.9165\pm 0.2820$ $0.7072\pm 0.1362$ $0.2321\pm 0.0847$ $0.9574\pm 0.1655$
Mole-BERT [37] $0.8624\pm 0.0732$ $1.2938\pm 0.3615$ $1.0529\pm 0.1536$ $1.4348\pm 0.1398$ $1.0440\pm 0.2569$ $0.9377\pm 0.1873$ $0.5190\pm 0.0645$ $1.4542\pm 0.3456$
ECFP [80] $1.1409\pm 0.2281$ $1.3721\pm 0.0703$ $1.2718\pm 0.1531$ $1.5446\pm 0.1133$ $1.2553\pm 0.0985$ $1.0439\pm 0.1244$ $1.4682\pm 0.2422$ $1.1965\pm 0.1850$
GraphMVP [38] $1.8865\pm 0.1303$ $2.8140\pm 0.3649$ $3.0341\pm 0.6418$ $2.7829\pm 0.1195$ $3.3965\pm 0.3437$ $3.0817\pm 0.2720$ $3.5723\pm 0.4337$ $3.0083\pm 0.2802$
NaFM $\mathbf{0.8243\pm 0.1960}$ $\mathbf{1.1227\pm 0.1604}$ $\mathbf{0.9239\pm 0.1721}$ $1.0802\pm 0.1506$ $\mathbf{0.6927\pm 0.2828}$ $\mathbf{0.6922\pm 0.1326}$ $\mathbf{0.2265\pm 0.0921}$ $\mathbf{0.9286\pm 0.1342}$